An investment of $12,000 is divided into two accounts whose annual simple interest rates are 9% and 11% How much has been deposited into each account if the interest earned by each account is the same?
Answer by Emsley(35)

Lrt amounts be x and y So x+y=12000 and 0.09x=0.11y so x=0.11y/0.09
Sub in other equation and get (0.11y/0.09)+y=12000
Multiply through by 0.09 and get 0.11y+0.09y= 1080 so 0.20y=1080 so y= 5400 and therefore x= 6600